# FedSDAF: Leveraging Source Domain Awareness for Enhanced Federated Domain Generalization

### v1

@@ 4.8. Visualization
We present several visualization results to provide an intuitive understanding of the advantages of FedSDAF. In Fig. 6, we apply the t-SNE algorithm (van der Maaten and Hinton, 2008) to visualize the features of target samples extracted by ViT-B/16, Domain Invariant Adapter, Domain Aware Adapter, and FedSDAF (KIA), respectively, with Sketch designated as the target domain on PACS. The t-SNE visualization reveals critical insights into FedSDAF’s feature discrimination capability within the challenging Sketch domain. While ViT-B/16 (Fig.6d) exhibits dispersed feature distribution with substantial inter-class overlap, both individual adapters demonstrate distinct adaptation strategies aligned with our design motivation: The Domain Invariant Adapter (Fig.6a) produces elliptical feature clusters with overlapping boundaries, reflecting its emphasis on cross-domain alignment at the expense of category discrimination. Conversely, the Domain Aware Adapter (Fig.6b) generates compact but partially misaligned clusters, indicative of its focus on capturing source-specific discriminative features while neglecting domain shift mitigation. FedSDAF (Fig.6c) fundamentally overcomes these limitations through hierarchical feature disentanglement. The isolated red cluster in the lower-left quadrant manifests the domain-aware component’s success in preserving source-discriminative patterns, while the uniform spacing between blue and purple clusters in the upper region verifies the domain-invariant module’s effectiveness in establishing cross-domain consistency. This dual mechanism achieves geometric synergy - the domain-aware adapter actively pushes class clusters apart by amplifying source-specific discriminators, while the domain-invariant adapter passively pulls features toward domain-agnostic decision boundaries. Such collaborative dynamics explain why FedSDAF surpasses both the base model and standalone adapters, as visualized by the sharpened cluster margins and reduced overlap density compared to other configurations. These visual patterns collectively validate our core motivation: Source-aware feature refinement requires simultaneous discriminative pattern preservation (through domain-aware adaptation) and cross-domain topology alignment (via invariant feature projection), which FedSDAF achieves through its dual-branch architecture.
5. Conclusion
In conclusion, the Federated Source Domain Awareness Framework (FedSDAF) represents a significant advancement in federated domain generalization by effectively leveraging source domain-aware features. This core contribution enhances model adaptation in environments characterized by data heterogeneity and privacy constraints. The dual adapter architecture, consisting of the Domain-Invariant Adapter and the Domain-Aware Adapter, is essential for maintaining stability while integrating domain-specific knowledge. This integration notably improves the model’s capacity to generalize across unknown target domains, addressing a key challenge faced by traditional federated learning methods.
However, certain limitations persist. The current study primarily focuses on specific domain datasets, which may not encompass the full variability encountered in real-world applications. Additionally, in this study, we transfer knowledge between the server and clients using logits of a public dataset, motivated by efficiency and privacy considerations. Although empirical evidence suggests that sharing logits of public datasets between the server and clients is more privacy-preserving than sharing model gradients or parameters (Fan et al., 2025; Liu et al., 2021c), there is no theoretical guarantee that this approach does not compromise the privacy of clients’ sensitive data. This issue warrants further investigation to ensure that privacy concerns are adequately addressed in future iterations of the framework.
Appendix A Research Methods
A.1. Part One:Recap: Generalization Adjustment Strategy
In order to mitigate the risk of extracted domain-invariant features being overly biased towards a specific source domain, we introduce a Generalization Adjustment (GA) strategy(Zhang et al., 2023c) during the aggregation phase. This section provides a formal description of the novel GA strategy designed for federated domain generalization. The GA mechanism incorporates a dynamic weight adjustment module on the server side, aimed at optimizing the global model’s generalization capability. The complete workflow encompasses four key phases.
Initially, in the Initialization Phase, The server initializes the global model parameters $\theta^{0}$ and sets uniform aggregation weights $a^{0}=(1/M,\ldots,1/M)$ for $M$ clients. Each client maintains a local copy of the model $\theta_{i}^{0}=\theta^{0}$ with domain-specific data $\widehat{D}_{i}$.
During the communication round $r$, the first step involves measuring the generalization gap. Each client computes the domain-specific generalization gap prior to local training, quantified by the equation:
(9) $\widehat{G}_{D}^{i}(\theta^{r})=\mathbb{E}^{\widehat{D}^{i}}(\theta^{r})-\mathbb{E}^{\widehat{D}^{i}}(\theta_{i}^{r-1})$
This equation captures the performance discrepancy between the current global model and the previous local optimum. Following this measurement, clients proceed with local model updates by performing E epochs of training using the algorithm to derive updated parameters
In the subsequent Server-side Weight Adjustment phase, the server undertakes two essential operations. The first is the momentum-based weight update, formulated as:
(10) $a_{i}^{r^{\prime}}=\left(\widehat{G}_{D}^{i}(\theta^{r})-\mu\right)\cdot\frac{d^{r}}{\max_{j}(\widehat{G}_{D}^{j}(\theta^{r})-\mu)}+a_{i}^{r-1}$
where $\mu$ denotes the average gaps and $d^{r}=(1-r/R)\cdot d$ implements a linear decay to control the update magnitude. Subsequently, the server conducts a generalization-aware aggregation, computed as:
(11) $\theta^{r+1}=\sum_{i=1}^{M}a_{i}^{r}\cdot\theta_{i}^{r^{\prime}}$
This reweighted aggregation approach supersedes the static weighting scheme employed by FedAvg.
Finally, in the Model Distribution phase, the refined global model $\theta^{r+1}$ is broadcast to all clients for subsequent rounds, thereby completing the closed-loop optimization cycle.
A.2. Part Two: Algorithm 1
As depicted in Algorithm 1 On the client side, each client initializes its local copy $\hat{A}_{di}^{k,r}$ corresponding to the received global domain-invariant adapter $A_{di}^{r}$. The client subsequently optimizes its local domain-aware adapter $A_{dw}^{k,r}$ and the local replica $\hat{A}_{di}^{k,r}$ through the bidirectional knowledge distillation mechanism. During this optimization, the client computes its training error $\hat{E}_{D_{k}}^{k,r}(A_{di}^{k,r})$ and, after each round, uploads the generalization gap $G_{k}^{r}$ (if $r>0$) along with the updated domain-invariant adapters $A_{di}^{k,r}$ back to the server. This systematic updating process enables the FedSDAF method to effectively leverage source domain knowledge, thereby enhancing the model’s generalization performance in unseen target domains.
Input: $K$: Number of clients
$\{D_{k}\}_{k=1}^{K}$: Client datasets
$R$: Communication rounds
Output: $\{A_{dw}^{k,R}\}_{k=1}^{K}$: Well-trained client models
$A_{di}^{R}$: Global domain-invariant adapter
1
2Server Execution:
3 Initialize $A_{di}^{0}$, $A_{dw}^{k,0}\leftarrow$ random initialization;
4 $a_{k}^{0}\leftarrow\frac{1}{M},\ \forall k\in\{1,\ldots,K\}$;
5
6for $r\leftarrow 0$ to $R-1$ do
7       if $r>0$ then
8             Collect $\{G_{k}^{r},A_{di}^{k,r}\}_{k=1}^{K}$;
9             $a_{k}^{r}\leftarrow\textsc{GA-Algorithm}(G_{k}^{r})$;
10             $A_{di}^{r+1}\leftarrow\sum_{k=1}^{K}a_{k}^{r}A_{di}^{k,r}$;
11
12       else
13             $A_{di}^{r+1}\leftarrow A_{di}^{0}$;
14
15       end if
16      Broadcast $A_{di}^{r+1}$;
17
18 end for
19
20Client $k$ Execution (Parallel):
21 for $r\leftarrow 0$ to $R-1$ do
22       Receive $A_{di}^{r}$;
23       $\hat{A}_{di}^{k,r}\leftarrow A_{di}^{r}$;
24       while not converged do
25             $\min_{A_{dw}^{k,r},\hat{A}_{di}^{k,r}}\mathcal{L}_{\text{distill}}(A_{dw}^{k,r},\hat{A}_{di}^{k,r})$;
26
27       end while
28      Compute $\hat{E}_{D_{k}}^{k,r}(A_{di}^{k,r})$;
29       if $r>0$ then
30             $G_{k}^{r}\leftarrow\hat{E}_{D_{k}}^{k,r}-\hat{E}_{D_{k}}^{k,r-1}$;
31
32       end if
33      Send $(G_{k}^{r},A_{di}^{k,r})$;
34
35 end for
Algorithm 1 Federated Domain Adaptation with Bidirectional Distillation (FedSDAF)

### v3

@@ B.2 Visualization
We present several visualization results to provide an intuitive understanding of the advantages of FedSDAF. In Fig. 6, we apply the t-SNE algorithm to visualize the features of target samples extracted by ViT-B/16, Domain Invariant Adapter, Domain Aware Adapter, and FedSDAF (KIA), respectively, with Sketch designated as the target domain on PACS. The t-SNE visualization reveals critical insights into FedSDAF’s feature discrimination capability within the challenging Sketch domain. While ViT-B/16 (Fig.6d) exhibits dispersed feature distribution with substantial inter-class overlap, both individual adapters demonstrate distinct adaptation strategies aligned with our design motivation: The Domain Invariant Adapter (Fig.6a) produces elliptical feature clusters with overlapping boundaries, reflecting its emphasis on cross-domain alignment at the expense of category discrimination. Conversely, the Domain Aware Adapter (Fig.6b) generates compact but partially misaligned clusters, indicative of its focus on capturing source-specific discriminative features while neglecting domain shift mitigation. FedSDAF (Fig.6c) fundamentally overcomes these limitations through hierarchical feature disentanglement. The isolated red cluster in the lower-left quadrant manifests the domain-aware component’s success in preserving source-discriminative patterns, while the uniform spacing between blue and purple clusters in the upper region verifies the domain-invariant module’s effectiveness in establishing cross-domain consistency. This dual mechanism achieves geometric synergy - the domain-aware adapter actively pushes class clusters apart by amplifying source-specific discriminators, while the domain-invariant adapter passively pulls features toward domain-agnostic decision boundaries. Such collaborative dynamics explain why FedSDAF surpasses both the base model and standalone adapters, as visualized by the sharpened cluster margins and reduced overlap density compared to other configurations. These visual patterns collectively validate our core motivation: Source-aware feature refinement requires simultaneous discriminative pattern preservation (through domain-aware adaptation) and cross-domain topology alignment (via invariant feature projection), which FedSDAF achieves through its dual-branch architecture.
B.3 Scalability with Increasing Client Pool Size
A critical aspect of any federated learning framework is its ability to scale effectively as the number of participating clients grows. To assess the scalability of FedSDAF, we conducted an experiment on the PACS dataset by varying the number of clients ($K$) among which the source domains are partitioned. We evaluated the framework’s performance with client pool sizes of $K=4,20,30,$ and $50$.
The results, presented in Figure 7, demonstrate that FedSDAF exhibits robust scalability. As the number of clients increases, there is a slight and graceful degradation in accuracy across all domains. This trend is expected in federated settings, as a larger client pool typically introduces greater data heterogeneity and makes achieving a global consensus more challenging. However, the performance decline is modest, even when scaling to 50 clients. For instance, on the most challenging Sketch (S) domain, the accuracy remains high, showcasing the framework’s ability to maintain its effectiveness. This highlights that FedSDAF’s core mechanisms, particularly the synergistic learning between the domain-aware and domain-invariant adapters, are resilient to increased decentralization, making the framework highly suitable for large-scale, real-world deployments.
Figure 7: Performance of FedSDAF on the PACS dataset with a varying number of clients ($K$). The results show a graceful and minor decrease in accuracy as the client pool size increases, demonstrating the framework’s strong scalability.
B.4 Analysis of Adapter Fusion Strategy
Our framework integrates the outputs of the Domain-Aware ($A_{dw}$) and Domain-Invariant ($A_{di}$) adapters using a simple summation with equal weights, as described in Equation (2): $h^{\prime}=h+\frac{1}{2}\hat{A}_{dw}(h)+\frac{1}{2}\hat{A}_{di}(h)$. To validate this design choice, we performed an ablation study to analyze the impact of different fusion strategies on the PACS dataset. We explored two alternatives: (1) modifying the relative weighting of the adapter outputs, and (2) replacing the summation with a concatenation operation.
The results are illustrated in Figure 8. The ’Weight’ parameter in the legend corresponds to the balance between the two adapters, where ’Weight=1’ represents our original equal-weighting scheme. The analysis reveals several key insights:
  • Robustness to Weighting: The model achieves the highest performance with ’Weight=1’ (our default) and ’Weight=1.5’, indicating that a balanced contribution is optimal. Performance degrades significantly when the weight is too low (e.g., ’Weight=0.1’), confirming that a substantial signal from both adapters is crucial for success.
  • Effectiveness of Concatenation: The ’Concat’ strategy, where the feature vectors from both adapters are concatenated, also yields excellent results. Its performance is highly competitive with the optimal weighted-sum approach across all domains.
This study confirms that our default equal-weighting scheme is a simple, robust, and effective choice. Furthermore, the strong performance of the concatenation alternative validates the core hypothesis of FedSDAF: the explicit disentanglement and subsequent reintegration of local expertise and global consensus are the primary drivers of performance, and the framework is not overly sensitive to the specific fusion mechanism employed.
Figure 8: Performance comparison of different adapter fusion strategies on the PACS dataset. Both the weighted-sum and concatenation methods prove effective, with a balanced weighting scheme performing optimally.
B.5 Ablation Study on MHSA via Dynamic Distillation Weighting
In our main experiments, the FedSDAF framework utilizes a Multi-Head Self-Attention (MHSA) mechanism alongside a fixed weight for the bidirectional knowledge distillation (BKD) loss. To validate the unique contribution of the MHSA component, we conducted an ablation study where we removed it and attempted to compensate by introducing a more sophisticated, dynamic weighting strategy for the distillation loss. Specifically, instead of a fixed $\alpha$, we employed an exponentially increasing weight governed by the formula:
$\alpha_{t}=\alpha_{\text{max}}\times\left(1-e^{-\frac{t}{\tau}}\right)$ (7)
where $t$ is the current epoch, $\alpha_{\text{max}}$ is the maximum weight, and $\tau$ controls the rate of increase. The goal was to see if an optimally tuned, dynamic distillation process could replicate the performance gains from MHSA.
We performed a grid search for $\alpha_{\text{max}}$ and $\tau$ on the PACS dataset for this MHSA-free variant. The results are shown in Figure 9.
Figure 9: Heatmap of average accuracy for a FedSDAF variant without MHSA, using a dynamic distillation weight $\alpha_{t}$. The analysis explores different values for $\alpha_{\text{max}}$ and $\tau$ on the PACS dataset.
The analysis of this experiment leads to a clear conclusion. The heatmap reveals that the optimal performance for the MHSA-free variant is achieved with a high $\alpha_{\text{max}}$ (2.0) and a moderately fast rate of increase ($\tau=100$), yielding a peak average accuracy of approximately 97.08%. While this result confirms that a carefully tuned dynamic distillation schedule can enhance performance, it crucially represents the performance ceiling for this ablated model configuration. Notably, this peak accuracy remains significantly below the performance of our full FedSDAF model, which incorporates the MHSA mechanism (as detailed in Table 5). This performance deficit underscores that the contribution of MHSA is not merely a matter of balancing feature fusion, which could be emulated by adjusting loss weights. Instead, it provides a distinct architectural advantage, likely by capturing complex inter-feature relationships that are essential for robust generalization. Therefore, this ablation study validates that while a dynamic distillation weight is a viable optimization strategy, it is not a substitute for the fundamental benefits conferred by the MHSA component, thereby justifying its integral role in our proposed framework.